# MACI: Multi-Agent Collaborative Intelligence for Robust Reasoning and Temporal Planning

### v1

@@ Abstract
Artificial intelligence requires deliberate reasoning, temporal awareness, and effective constraint management—capabilities beyond the pattern-matching strengths of LLMs. LLMs struggle with planning tasks because of their reliance on associative reasoning, inability to self-verify, and inconsistent constraint awareness. We propose Multi-Agent Collaborative Intelligence (MACI), a framework centered on a meta-planner (MP) that orchestrates multiple agents to generate planner templates that define roles and constraints. These planners produce actionable workflows of role nodes and dependency constraints, enabling advanced temporal reasoning and adaptability.
MACI’s three-tier architecture includes a meta-planning module for planner construction, common agents for general reasoning, and specialized agents for domain expertise. By decoupling planning from validation, it overcomes key LLM limitations. Evaluations demonstrate MACI’s effective constraint satisfaction, conflict detection, and reasoning, positioning it as a robust solution for complex reasoning and planning tasks.
@@ 1 Introduction
The development of artificial general intelligence requires a shift from pattern matching to deliberate analytical reasoning. Large Language Models (LLMs) excel at pattern matching but lack deliberative reasoning required for planning tasks (Kahneman, 2011). These capabilities become crucial for tasks requiring sustained attention, awareness of constraints, and reasoning about future states.
@@ 1.1 Limitations of LLMs in Planning
Our evaluation reveals three critical limitations in current LLMs that prevent effective planning.
Metacognitive Limitations in Self-verification
LLM limitations in self-verification extend beyond Gödel’s incompleteness theorems for formal systems (Gödel, 1967). Though formal systems cannot prove their own consistency within their framework, LLMs face greater challenges due to their probabilistic nature and lack of logical foundations (detailed arguments in Section 2). This limitation provides theoretical justification for distributed validation, where specialized agents perform independent verification of planning processes.
Attention-based Processing Bias
In complex planning scenarios, LLMs demonstrate a form of cognitive tunneling where recent context dominates earlier-established constraints. For example, when planning a multi-stop journey, an LLM might focus on optimizing the final leg’s arrival time while losing sight of earlier constraints about vehicle availability or required rest periods. This attention bias leads to locally optimized but globally invalid solutions, undermining the overall feasibility of the plan. The inability to maintain consistent awareness of all constraints throughout the planning process represents a fundamental limitation in current architectures.
Common Sense Knowledge Integration
LLMs often overlook practical constraints that humans naturally consider. In travel planning, they might ignore factors such as airport security wait time or luggage claim delays. These omissions are the result of an inadequate integration of common sense knowledge into the planning process. Although humans automatically incorporate such practical considerations, LLMs require explicit specification of these constraints, often leading to incomplete or impractical plans.
Table 1: Comparison of Validation Limitations and Solutions
Gödel’s Incomp. Theorems LLM Validation Limitations MACI’s Solution
A formal system cannot prove its own consistency LLMs cannot reliably validate their own outputs due to probabilistic foundations MACI uses external validation agents to bypass self-referential loops
Consistency proofs require a “higher” system Validation requires external frameworks (e.g., logic engines, human oversight) Distributed validation acts as the “higher" system for LLM outputs
Self-reference leads to paradox or incompleteness LLMs hallucinate or contradict themselves when critiquing their own outputs Meta-planning separates planning and validation into distinct agent roles
@@ 1.2 The MACI Framework
To overcome the inherent limitations of single-model approaches in planning tasks, we introduce the Multi-Agent Collaborative Intelligence (MACI) framework. At the core of MACI lies the meta-planner (MP), a high-level orchestration mechanism that collaborates with multiple agents to construct a specialized planner for temporal reasoning and scheduling. By enabling agents to work collectively at the meta-level, MACI ensures robust and adaptive planning across complex domains.
Meta-Planning Module
The meta-planning module serves as the cornerstone of MACI, analyzing task requirements and dynamically generating a planner specifically tailored to address complex challenges. This planner, in turn, produces a workflow composed of nodes representing roles (e.g., cook, driver, supervisor) and dependencies representing constraints (e.g., temporal, spatial, resource). Unlike traditional multi-agent systems, MACI’s meta-planner orchestrates a dynamic network of agents, validates their interactions, and enforces protocols to ensure comprehensive planning and execution. This layered approach—planning to generate a planner that produces actionable workflows—positions MACI as a forward-looking solution to real-world planning problems.
Common Agents
@@
Common agents form the foundation of MACI’s general-purpose planning capabilities. They address critical tasks such as practical reasoning, constraint validation, and robustness analysis. For example, the Common Sense Integration agent identifies implicit constraints, the Constraint Validation agent ensures solution feasibility, and the Plan Evaluation agent measures performance metrics. These agents collectively maintain both explicit and implicit constraints while ensuring compliance and safety across domains.
Specialized Agents
Specialized agents extend MACI’s capabilities by addressing domain-specific requirements. Instead of overburdening a single planner, MACI deploys focused agents for tasks such as travel scheduling, financial risk assessment, or logistics optimization. These agents integrate domain expertise with the foundational work of common agents, ensuring precise and effective solutions for specialized planning challenges.
1.3 Contributions
This work introduces MACI as a novel approach to collaborative planning. Preliminary evaluations in travel planning scenarios demonstrate MACI’s superiority in constraint satisfaction, conflict detection, and practical reasoning compared to single-LLM approaches. Key contributions of this paper include:
  1. 1.
    Identifying limitations of single-LLM approaches through systematic experimentation.
  2. 2.
    Developing a three-tier architecture for distributed planning and validation.
  3. 3.
    Designing MACI with common and specialized agents for constraint management and reasoning.
  4. 4.
    Demonstrating MACI’s effectiveness in travel planning scenarios, achieving significant improvements over single-LLM systems.
The remainder of this paper is organized into five sections. Section 2 reviews related MAS frameworks and explains why they fall short for the task at hand. Section 3 examines the limitations of LLMs through case studies. Section 4 details the specifications of the MACI framework. Section 5 evaluates MACI’s performance in comparison to state-of-the-art LLMs. Finally, Section 6 discusses implications for future research and alternative perspectives.
@@ 2 Related Work
The development of MACI builds on theoretical insights from formal systems and addresses limitations of current multi-agent architectures. Gödel’s second incompleteness theorem (Kennedy, 2008; Gödel, 1967) established that no consistent formal system can prove its own consistency. We extend this insight to LLMs, which operate on probabilistic rather than axiomatic foundations, making them inherently incapable of reliable self-validation. Table 1 illustrates how MACI’s distributed validation architecture addresses these fundamental limitations.
@@
Current multi-agent systems (MAS) primarily serve as integration platforms for coordinating multiple LLMs. Notable frameworks include Microsoft’s AutoGen (Wu et al., 2024), the Multi-LLM Agent Debate Framework (Du et al., 2023; Chang, 2023, 2024a, 2024b), LangGraph and CrewAI (LangChain AI, 2024; Moura, 2024), XAgent (Xia et al., 2023), and CAMEL (Li et al., 2023). Although these frameworks excel at agent coordination, they focus on task distribution rather than comprehensive constraint management required for complex planning.
MACI’s meta-planning module (Section 4.1) addresses these limitations by acting as a metasystem that coordinates specialized agents. By decoupling planning and validation functions, MACI creates a validation framework analogous to the higher-order system required to demonstrate formal consistency while maintaining awareness of interdependent constraints.
3 Case Study: Can LLMs Handle Planning
This section begins by presenting various types of planning, including strategic approaches and tactical methodologies. We then present experimental results on a scheduling problem to illustrate how current LLMs fall short. Based on our diagnoses, we propose remedies in Section 4.
Planning Types
Planning can be classified into seven types. The first three are considered more strategic and the rest are tactic methodologies.
  1. 1.
    Sequential planning: Creates time-ordered schedules meeting goals/constraints via formal temporal representations (Allen & Hayes, 1989)
  2. 2.
    Proactive planning: Anticipates future scenarios to preemptively address issues through deliberate foresight (Cox & Veloso, 1998)
  3. 3.
    Retrospective planning: Improves future plans by analyzing past experiences via case-based reasoning (Kolodner, 1993)
  4. 4.
    Adaptive planning: Adjusts plans to changing conditions while preserving core constraints (Hammond, 1990)
  5. 5.
    Reactive planning: Prioritizes immediate actions over long-term goals in dynamic, uncertain environments (Georgeff & Lansky, 1987)
  6. 6.
    Predictive planning: Uses data-driven models to forecast scenarios and optimize decisions (Kushmerick et al., 1995)
@@ 3 Case Study: Can LLMs Handle Planning
Using this problem, three key types of planning conducted by GPT4o and DeepSeek are examined: sequential, proactive/reactive, and retrospective.
@@
3:00pm James arrives at airport to pick up Emily J
@@
3:15pm Michael at Grandma’s home. M
@@ 3.1 Study #1 Sequential Planning
5:30pm Michael arrives home with Grandma M
@@
Analysis (Appendix A) links both errors to flawed reasoning in constraint interpretation.
@@
3.2 Study #2 Proactive and Reactive Planning
@@
3:00pm Sarah departs to pick up Grandma Sarah
@@
3.3 Study #3 Retrospective Planning
Retrospective planning improves future strategies by learning from past decisions. Both DeepSeek and GPT-4o lacked flexibility even after error correction: DeepSeek relied on rigid templates, and GPT-4o’s initial assignment (Sarah homebound in Table 3) limited later options.
Missed Opportunities During James’ 3-Hour Delay:
  • Sarah could have picked up Grandma earlier, freeing her for contingencies like Emily’s transport.
  • Michael could have left NYC at 10:00 AM to reach Emily by 2:30 PM.
In Section 4, we show how MACI’s meta-planner effectively revises plans by debugging errors and adapting to dynamic constraints.
@@ 4 MACI Framework Specification
MACI implements a three-layer architecture to address current LLM limitations: metacognitive constraints, attention bias, and gaps in common-sense reasoning. The top two layers represent key innovations, while the bottom layer leverages existing multi-agent system (MAS) infrastructure.
4.1 Three-Layer Architecture
Meta-Planner: The Planning Generator
@@ 4.1 Three-Layer Architecture
$\mathcal{MP}:(\mathcal{O},\mathcal{C_{E}})\rightarrow\mathbf{W},$
where $\mathbf{W}$ represents a planning system composed of specialized, coordinated agents. Similar to how a compiler generator produces compilers from specifications, $\mathcal{MP}$ constructs agent networks from task requirements. It analyzes objectives, identifies required roles and dependencies, selects appropriate agents, and establishes interaction protocols. This produces a complete workflow template defining both the planning state space and the agent coordination mechanisms needed to solve the given task.
Agent Repository Layer
This layer maintains a distributed collection of planning agents, each designed with a restricted context window and specialized interface. By dividing cognitive tasks among agents, this architecture ensures a complete knowledge of the constraints without overwhelming individual components.
System Infrastructure Layer
Built on open-source MAS frameworks, the infrastructure layer supports essential operations such as agent registration, message routing, resource allocation, and deployment scaling.
4.2 Agent Repository Layer Design
4.2.1 Agent Registration and Search Protocol
Agents register in the repository with five attributes: input/output protocol ($P$), type ($t$), capability vector ($c$), context window size ($w\leq 1k$ tokens) and user rating ($r$). These attributes enable $\mathcal{MP}$ to match task requirements to agent capabilities efficiently. $\mathcal{MP}$ searches for agents by: 1) matching task requirements to capability descriptions ($c$), 2) filtering by agent type ($t$) and user ratings ($r$), and 3) verifying protocol compatibility ($P$).
@@ 4.2.2 State Space and Agent Design
General State Space
Table 5: State Space Dimensions and Components
Dimension Core Components Example States
Who
• Actor ID
• Current Roles
• Role Constraints
• Role Changes
• Driver vs Passenger
• Supervisor vs Worker
• Buyer vs Seller
• Multiple role conflicts
Where
• Current Position
• Target Location
• Transition Points
• Access Rules
• Physical locations
• Virtual positions
• State transitions
• Boundary constraints
When
• Time Points
• Duration
• Deadlines
• Probability
• Event timestamps
• Process duration
• Completion times
• Delay likelihood
What
• Methods
• Requirements
• Constraints
• Costs
• Tools/Vehicles
• Tickets/Permits
• Capacity limits
• Time-money trade-offs
Why
• Rationale
• Dependencies
• Risks
• Alternatives
• Decision basis
• Causal chains
• Failure modes
• Backup plans
Table 5 presents five dimensions of state space, Who, Where, When, What, and Why, that form the basis of planning tasks. These dimensions are general enough to apply across domains, with travel planning serving as an illustrative example.

### v2

# MACI: Multi-Agent Collaborative Intelligence for Adaptive Reasoning and Temporal Planning

@@ Abstract
Artificial intelligence requires deliberate reasoning, temporal awareness, and effective constraint management—capabilities traditional LLMs often lack due to their reliance on pattern matching, limited self-verification, and inconsistent constraint handling. We introduce Multi-Agent Collaborative Intelligence (MACI), a framework comprising three key components: 1) a meta-planner (MP) that identifies, formulates, and refines all roles and constraints of a task (e.g., wedding planning) while generating a dependency graph, with common-sense augmentation to ensure realistic and practical constraints; 2) a collection of agents to facilitate planning and address task-specific requirements; and 3) a run-time monitor that manages plan adjustments as needed. By decoupling planning from validation, maintaining minimal agent context, and integrating common-sense reasoning, MACI overcomes the aforementioned limitations and demonstrates robust performance in two scheduling problems.
@@ 1 Introduction
Advancing artificial intelligence requires capabilities beyond pattern matching. To tackle complex real-world tasks, AI must exhibit deliberate reasoning, temporal awareness, and effective constraint management. While Large Language Models (LLMs) excel at pattern recognition, they face significant challenges in planning tasks that demand sustained attention, comprehensive constraint awareness, and reasoning across both past and future temporal states (Kahneman, 2011).
@@ 1.1 Limitations of LLMs in Planning
LLMs reveal three limitations that fundamentally undermine their effectiveness in complex planning scenarios:
1. Lack of Self-Verification. LLMs struggle with validating their own output, a problem that extends beyond Gödel’s incompleteness theorems for formal systems (Gödel, 1967). Their probabilistic nature and lack of logical foundations create significant barriers to self-assessment. This intrinsic limitation means LLMs cannot reliably detect errors or inconsistencies in their generated content, necessitating external mechanisms to validate and refine their output.
2. Attention Bias and Constraint Drift. In complex scenarios, LLMs demonstrate a critical cognitive limitation known as cognitive tunneling. This phenomenon occurs when recently provided context dominates and progressively erodes earlier-established constraints. When planning a multi-leg journey, for instance, an LLM might optimize the final travel segment while completely neglecting crucial earlier constraints such as vehicle availability or required rest periods. This bias toward local optimization fundamentally undermines the global feasibility of generated plans.
3. Lack of Common Sense Integration. LLMs often overlook practical constraints that humans intuitively consider. This deficiency becomes particularly evident in domains that require real-world experience and understanding. In travel planning, an LLM might generate a route without accounting for airport security processing times. In logistics, it may create schedules that ignore resource availability and preparation windows. Without explicit, granular specifications, these models produce plans that appear superficially coherent but remain impractical.
@@ 1.2 The MACI Framework
To address these limitations, we propose Multi-Agent Collaborative Intelligence (MACI), a framework designed to enhance reasoning and planning through a multi-component architecture. MACI introduces three core components:
1. Meta-Planner (MP). The meta-planner serves as the central orchestration mechanism in MACI. It analyzes task requirements, identifies roles and constraints, and dynamically generates a dependency graph (or workflow template) tailored to the task. This template includes actionable workflows with nodes representing roles (e.g., cook, driver, supervisor) and edges representing dependencies (e.g., temporal, spatial, or resource constraints). The incorporation of common sense augmentation into the metaplanner ensures that the generated plans are realistic, comprehensive, and aligned with practical constraints.
2. Common and Task-specific Agents. MACI employs two types of agents to execute the generated plans:
@@
    Common Agents: These agents handle general-purpose tasks, including constraint validation, practical reasoning, and performance evaluation. For instance, a Common Sense Integration agent identifies implicit constraints, while a Constraint Validation agent ensures feasibility and compliance with the task’s requirements.
    Task-specific Agents: These agents cater to domain-specific requirements, including task-dependent data and knowledge augmentation, selection of the most effective planning algorithms, safety and ethics assessment, and emergency response optimization. By integrating domain expertise, they extend the capabilities of common agents, enabling MACI to address specialized planning challenges with precision and adaptability.
3. Run-Time Monitor. The run-time monitor handles real-time adjustments to the static plan in response to unexpected changes, such as resource delays, environmental disruptions, or evolving task requirements. This component ensures adaptability and robustness by:
    Monitoring plan execution to detect deviations.
    Activating emergency agents to revise dependencies, reassign roles, or dynamically adjust constraints.
    Communicating updates to affected agents to maintain coherence throughout the workflow.
1.3 Summary: How MACI Addresses LLM Limitations
With its multi-component architecture, MACI directly addresses the three critical limitations of LLMs in planning:
1. Lack of Self-Verification. MACI separates planning from validation, employing independent agents for validation. These agents operate without shared memory or interference, ensuring external verification of outputs and mitigating the risks of self-referential errors.
2. Attention Bias and Constraint Drift. MACI avoids relying on a single LLM to execute complex, multi-step reasoning sequentially. Instead, it utilizes small collaborative agents that enjoy two key benefits: independence and well-defined input/output protocols (ensuring specificity and quality) for specific tasks. These agents operate within restricted context windows of e.g., 1k tokens, which physically limits attention bias and ensures that earlier constraints are not overridden by recent context. By logically scoping problems and physically constraining context, MACI preserves global feasibility and mitigates cognitive tunneling.
Lack of Common Sense Integration. MACI incorporates a Common Sense Integration Agent and other specialized agents to identify implicit constraints and augment plans with practical, domain-specific knowledge. This ensures that generated plans are realistic, comprehensive, and aligned with real-world conditions.
Through its innovative architecture, MACI overcomes the inherent limitations of LLMs, enhancing their capacity for deliberate reasoning and planning. In subsequent sections, we demonstrate MACI’s effectiveness through evaluations in complex scenarios, such as the Traveling Salesman Problem (TSP) and a multi-layered dinner planning task.
@@ 2 Related Work
The development of MACI builds on theoretical insights from formal systems and addresses limitations of current multi-agent architectures. Gödel’s second incompleteness theorem (Kennedy, 2008; Gödel, 1967) established that no consistent formal system can prove its own consistency. This principle extends to LLMs, which rely on probabilistic rather than axiomatic foundations, making them inherently incapable of reliable self-validation. To address this, MACI employs a distributed validation architecture, where independent agents validate externally the output, bypassing the self-referential loops that may lead to inconsistencies.
In formal systems, consistency proofs require a “higher-order” system. Analogously, MACI provides a validation framework that operates as a higher-order metasystem for LLM output. By decoupling planning from validation, MACI mirrors the separation needed in formal systems, where validation is performed independently to avoid conflicts and errors.
Moreover, MACI advances the state of the art in multi-agent systems by addressing challenges that existing frameworks have not fully resolved.
@@
Current multi-agent systems (MAS) primarily function as integration platforms for coordinating multiple LLMs. Notable frameworks include Microsoft’s AutoGen (Wu et al., 2024), the Multi-LLM Agent Debate Framework (Du et al., 2023; Chang, 2023, 2024a, 2024b), LangGraph and CrewAI (LangChain AI, 2024; Moura, 2024), XAgent (Xia et al., 2023), and CAMEL (Li et al., 2023). While these frameworks excel in agent coordination, they prioritize task distribution over the comprehensive constraint management necessary for complex planning.
MACI bridges this gap by integrating a meta-planning module with independent agents that validate constraints, enabling robust and adaptable solutions in dynamic real-world scenarios. The meta-planner constructs task-specific dependency graphs that encode inter-agent constraints, ensuring precise role allocation while maintaining global feasibility. Meanwhile, validation agents, operating independently of the planning process, monitor for errors and inconsistencies stemming from probabilistic output, ensuring alignment with task objectives. This separation of roles mitigates cognitive tunneling and enhances adaptability, allowing MACI to dynamically respond to real-time disruptions such as resource shortages or environmental changes.
By integrating these advanced mechanisms, MACI goes beyond existing MAS frameworks to provide a cohesive architecture for complex reasoning and planning. It ensures a high degree of scalability and robustness, making it suitable for applications ranging from logistical optimization to adaptive decision-making in uncertain environments.
3 Case Study: Illuminating LLM Limitations
Planning methodologies can be broadly categorized into sequential and reactive approaches. Sequential planning involves creating time-ordered schedules (Allen & Hayes, 1989), anticipating future scenarios (Cox & Veloso, 1998), and leveraging past experiences for improvement (Kolodner, 1993). Reactive planning, on the other hand, focuses on adapting to changing conditions (Hammond, 1990), prioritizing immediate actions in dynamic environments (Georgeff & Lansky, 1987), and using data-driven models to forecast scenarios (Kushmerick et al., 1995).
This section highlights the limitations of current LLMs in planning tasks through experiments in two contexts: a scheduling problem illustrating shortcomings in sequential planning, and a dynamic resource allocation problem revealing challenges in reactive planning. Based on these observations, we propose remedies in Section 4.
@@ 3 Case Study: Illuminating LLM Limitations
Using this problem, we examine sequential and reactive planning performed by GPT-4o and DeepSeek.
@@
3:00pm X James arrives at airport to pick up Emily J
@@
3:15pm X Michael at Grandma’s home. M
@@ 3.1 Study #1 Sequential Planning
5:30pm X Michael arrives home with Grandma M
@@
Analysis (with detailed execution in Appendix A) links both errors to flawed reasoning in constraint interpretation.
@@
3.2 Study #2 Reactive Planning
@@
3:00pm X Sarah departs to pick up Grandma Sarah
@@ 3.3 Summary of LLM Limitations in Planning
In Sections 4 and 5, we show how MACI’s meta-planner effectively revises plans by debugging errors and adapting to dynamic constraints.
@@ 4 MACI Framework Specification
MACI implements a three-component architecture to address current LLM limitations: metacognitive constraints, attention bias, and gaps in common-sense reasoning. Each component plays a distinct role in enabling robust and adaptable planning capabilities.
4.1 Three-Component Architecture
Meta-Planner Component
@@ 4.1 Three-Component Architecture
$\mathcal{MP}:(\mathcal{O},\mathcal{C_{E}})\rightarrow\mathbf{W},\vspace{-.05in}$
where $\mathbf{W}$ represents a planning system composed of specialized, coordinated agents. Similar to a compiler generator producing compilers from specifications, $\mathcal{MP}$ constructs agent networks from task requirements. It analyzes objectives, identifies required roles and dependencies, selects appropriate agents, and establishes interaction protocols. This produces a workflow template that defines the planning state space and the coordination mechanisms needed to solve the task.
Agent Repository Component
This component maintains a distributed collection of planning agents, each designed with a restricted context window and specialized interface. By dividing cognitive tasks among agents, the repository ensures a complete representation of constraints without overwhelming individual components. The meta-planner queries this repository to select agents for specific roles and dependencies based on task requirements.
System Infrastructure Component
Built on open-source multi-agent system (MAS) frameworks, the infrastructure component supports essential operations such as agent registration, message routing, resource allocation, and deployment scaling. This foundation provides the necessary runtime environment for executing and monitoring the generated workflows.
4.2 Agent Repository Design
The agent repository in MACI serves as a structured database, enabling efficient registration, retrieval, and matching of agents to task requirements. By categorizing agents into common agents and specialized agents, the repository supports both generalized functionality and domain-specific expertise, as outlined in Section 4.1.
4.2.1 Lightweight, Independent Agent Design
MACI avoids relying on a single LLM to execute complex, multi-step reasoning sequentially. Instead, it utilizes small, independent agents that adhere to strict efficiency and modularity principles. These agents operate with well-defined input/output protocols and are constrained to restricted context windows to mitigate attention bias and prevent earlier constraints from being overridden by recent context.
By scoping problems logically and constraining context physically, MACI ensures that each agent processes only the task-relevant information needed for its specific role. This design prevents cognitive tunneling, maintains global feasibility, and enhances robustness in dynamic environments.
4.2.2 Agent Registration and Specifications
Each agent is registered in the repository using a standardized protocol buffer that encodes the following attributes:
  • Input/output protocol ($P$): Defines the data format and expected interactions for seamless communication.
  • Agent type ($t$): Specifies whether the agent is common or specialized.
  • Capability vector ($c$): Encodes the agent’s functional capabilities, constraints, and operating conditions.
  • Context window size ($w$): Ensures that each agent operates within a restricted buffer ($w\leq 1k$ tokens) to prevent attention bias and excessive information retention.
  • Computational efficiency constraint ($e$): Agents are lightweight, avoiding unnecessary memory usage or processing delays.
  • User rating ($r$): Tracks historical performance evaluations to prioritize reliable agents during selection.
The meta-planner retrieves agents from the repository using a three-step matching process:
  1. 1.
    Task-to-capability matching: Filters agents based on their capability vector ($c$) and task-specific requirements.
  2. 2.
    Protocol verification: Ensures compatibility of input/output protocols ($P$) between selected agents to prevent communication errors.
  3. 3.
    Agent ranking: Ranks agents by their relevance, efficiency, and historical user rating ($r$) to select the optimal candidates.
This structured retrieval mechanism ensures that MACI efficiently scales to complex planning problems without requiring predefined agent hierarchies. By leveraging protocol buffers and a structured repository, MACI achieves both modularity and adaptability, allowing new agents to be introduced seamlessly while maintaining coherence across multi-agent interactions.
@@ 4.2.3 State Space and Agent Design
Tasks in MACI are modeled in a general five-dimensional state space to ensure comprehensive representation of constraints and dependencies. These dimensions include:
  1. 1.
    Who (Actors): Identifies roles, constraints, and transitions between agents or individuals.
  2. 2.
    Where (Location): Tracks physical or logical positions, transitions, and access rules.
  3. 3.
    When (Time): Captures temporal constraints such as deadlines, durations, and time points.
  4. 4.
    What (Resources): Manages resource availability, constraints, and associated costs.
  5. 5.
    Why (Logic): Encodes rationale, dependencies, and risk assessments for decision-making.
This structured state space allows the meta-planner to generate workflows that account for all relevant constraints and dependencies across diverse domains.
4.2.4 Agent Roles in State Space Management